# Algorithmic contiguity from low-degree conjecture and applications in correlated random graphs

### v3

@@ 1.4 Organization of this paper
The rest part of this paper is organized as follows: in Section 2 we state the precise framework of low-degree polynomials and how we relate it to the notion of algorithmic contiguity (see Theorem 2.4). In Section 3 we use this framework to deduce the hardness of partial matching in correlated Erdős-Rényi models and correlated SBMs, thus verifying Items (1) and (2) in Theorem 1.6 (see Theorem 3.2). In Section 4 we again use this framework to deduce the hardness of testing correlated SBMs against independent SBMs, thus verifying Item (3) in Theorem 1.6 (see Corollary 4.2). Several auxiliary results are moved to the appendix to ensure a smooth flow of presentation.
@@ 2 Low-degree framework and algorithmic contiguity
The low-degree polynomial framework first emerged from the works of [BHK+19, HS17, HKP+17, Hop18] and has since been refined and extended in various directions. It has found applications across a broad spectrum of problems, including detection tasks such as planted clique, planted dense subgraph, community detection, sparse PCA, and tensor PCA (see [HS17, HKP+17, Hop18, KWB22, SW22, DMW23+, BKW20, DKW+22, MW25, DDL23+, KMW24]), optimization problems like finding maximal independent sets in sparse random graphs [GJW20, Wein22], and constraint satisfaction problems such as random $k$-SAT [BH22]; see also the survey [KWB22]. Additionally, it is conjectured in [Hop18] that the failure of degree-$D$ polynomials implies the failure of all “robust” algorithms with running time $n^{\widetilde{O}(D)}$ (here $\widetilde{O}$ means having at most this order up to a $\operatorname{poly}\log n$ factor). In the remaining of this paper, we will focus on applying this framework in the context of high-dimensional hypothesis testing problems.
@@
We say an algorithm $\mathcal{A}$ that takes $\mathsf{Y}$ as input and outputs either 00 or $1$ achieves
  • strong detection, if the sum of type-I and type-II errors $\mathbb{Q}(\mathcal{A}(\mathsf{Y})=1)+\mathbb{P}(\mathcal{A}(\mathsf{Y})=0)$ tends to 00 as $n\to\infty$.
@@
${}\mathsf{Adv}_{\leq D}\Big{(}\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\Big{)}:=\sup_{f\in\mathcal{P}_{D}}\frac{\mathbb{E}_{\mathbb{P}}[f]}{\sqrt{\mathbb{E}_{\mathbb{Q}}[f^{2}]}}\,.$ (2.1)
@@
For “natural” high-dimensional hypothesis testing problems between $\mathbb{P}$ and $\mathbb{Q}$, the following statements hold:
@@
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=O(1)$ as $n\to\infty$, then there exists a constant $C$ such that there is no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves strong detection between $\mathbb{P}$ and $\mathbb{Q}$.
@@
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=1+o(1)$ as $n\to\infty$, then there exists a constant $C$ such that there is no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves weak detection between $\mathbb{P}$ and $\mathbb{Q}$.
Motivated by [Hop18, Hypothesis 2.1.5 and Conjecture 2.2.4] as well as the fact that low-degree polynomials capture the best known algorithms for a wide variety of statistical inference tasks, Conjecture 2.2 appears to hold for distributions of a specific form that frequently arises in high-dimensional statistics. For further discussion on what types of distributions are suitable for this framework, we refer readers to [Hop18, KWB22, Kun21, ZSWB22]. In addition, we point out that although in most applications (and in the statement of [Hop18, Hypothesis 2.1.5]) it is typically to take $\mathbb{Q}$ to be a “null” measure and $\mathbb{P}$ to be a “planted” measure (which makes (2.1) more tractable), several recent works [RSWY23, KVWX23] showed that this framework might also be applicable for many “planted-versus-planted” problems. Nevertheless, in this paper, we adopt a more conservative view and will explicitly indicate whenever $\mathbb{Q}$ is treated as a planted measure.
The framework in Conjecture 2.2 provides a useful tool for probing the computational feasibility of strong or weak detection. However, as discussed in Section 1.1, it turns out that the failure of strong detection is not enough in our cases, especially when we hope to perform some reductions between statistical models in a regime where weak (but not strong) detection is possible. Thus, in this regime, we aim to characterize a stronger framework that rules out all one-sided test. This motivates thus to propose the following notion of algorithmic contiguity.
@@
For “natural” high-dimensional hypothesis testing problems between $\mathbb{P}$ and $\mathbb{Q}$, we say an algorithm $\mathcal{A}$ that takes $\mathsf{Y}$ as input and outputs either 00 or $1$ is a $\mathbb{Q}$-based one-sided test, if
@@
The proof of Theorem 2.4 is incorporated in Section B.1 of the appendix. In the remaining part of this work we will also need a strengthening of Conjecture 2.2, as introduced in Section 1.1. As discussed in the beginning of this section, the low-degree conjecture asserts that (for certain testing problems) low-degree polynomials are at least as powerful as all algorithms of the corresponding runtime (where the correspondence is described in Conjecture 2.2). For our purposes, we introduce a natural refinement of the low-degree conjecture, which posits that low-degree polynomials perform at least as well as all algorithms of the corresponding runtime in terms of the value of the ratio on the right-hand side of (2.1).
Conjecture 2.5 (Revised low-degree conjecture).
For “natural” high-dimensional hypothesis testing problems between $\mathbb{P}$ and $\mathbb{Q}$, the following statements hold:
  1. (1)
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$ as $n\to\infty$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$, then there is a constant $C$ such that there is no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves strong detection between $\mathbb{P}$ and $\mathbb{Q}$.
  2. (2)
    If there is no algorithm with running time $n^{D}$ that achieves strong detection between $\mathbb{P}$ and $\mathbb{Q}$, then for all statistics $f=f(\mathsf{Y})$ that can be computed in running time $n^{D}$, there exists some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ (we allow $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ depend on $f$) such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$ such that $\mathsf{Adv}(f)=O(1)$, where
@@
    $\mathsf{Adv}(f)=\frac{\mathbb{E}_{\mathbb{P}^{\prime}}[f]}{\sqrt{\mathbb{E}_{\mathbb{Q}^{\prime}}[f^{2}]}}\,.$
  3. (3)
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=1+o(1)$ as $n\to\infty$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$, then there is a constant $C$ such that there is no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves weak detection between $\mathbb{P}$ and $\mathbb{Q}$.
  4. (4)
    If there is no algorithm with running time $n^{D}$ that achieves weak detection between $\mathbb{P}$ and $\mathbb{Q}$, then for all statistics $f=f(\mathsf{Y})$ that can be computed in running time $n^{D}$, there exists some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ (we allow $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ depend on $f$) such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$ such that $\mathsf{Adv}(f)=1+o(1)$).
Note that in Conjecture 2.5 we are allowed to replace $\mathbb{P},\mathbb{Q}$ with some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ that are statistically indistinguishable with $\mathbb{P},\mathbb{Q}$. This modification enables us to avoid some situations where the low-degree advantage explodes due to some “rare events” (see [BAH+22, DDL23+, DMW23+] for example). [^1] In fact, as pointed out by Alexander S. Wein (who learned the following example from Ansh Nagda) in personal communication with the author, such truncation appears necessary for the conjecture to hold. Consider the sparse-PCA problem where the observation $Y=\tfrac{\lambda vv^{\top}}{\|v\|^{2}}+W$ where $W$ is an $n*n$ GOE matrix and $v\in\mathbb{R}^{n}$ having coordinates i.i.d. drawn from the sparse Rademacher prior with parameter $\rho=o(1)$. In this case, when $10<\lambda<$ 1 we have $\mathbb{E}_{\mathbb{P}}[f]/\sqrt{\mathbb{E}_{\mathbb{Q}}[f^{2}]}=O(1)$ for all low-degree polynomials $f$ but $\mathbb{E}_{\mathbb{P}}[f]/\sqrt{\mathbb{E}_{\mathbb{Q}}[f^{2}]}=\omega(1)$ if $f$ is a thresholding function on $v_{0}^{\top}Yv_{0}$ where $v_{0}$ is a fixed sparse $\{0,1\}$-vector. However, our truncated conjecture still holds as we can take $\mathbb{P}^{\prime}$ to be the conditional measure of $\mathbb{P}$ given $v_{0}^{\top}Yv_{0}$ is small. This revised low-degree conjecture was first proposed in [MW23+] (in a slightly different manner) for a different purpose where they aimed to study the limits of precise error of all computationally feasible algorithms. However, we will show in the next two sections that this conjecture is also useful in performing reductions between different inference tasks.
@@ 3 Partial recovery in correlated random graphs
Given a sample $(A,B)$ from the law a pair of correlated random graphs in Definition 1.1 (we denote this law of $\mathbb{P}_{*}$). We say an algorithm $\mathcal{A}$ achieves strong partial matching, if it takes $(A,B)$ as input and outputs a family of estimators $\{h_{i,j}:1\leq i,j\leq n\}$ such that
@@
    $h_{i,1}+\ldots+h_{i,n}=1$ for all $1\leq i\leq n$ a.s. under $\mathbb{P}_{*}$;
@@
    There exists a fixed constant $\iota>0$ such that
    $\mathbb{P}_{*}\big{(}h_{1,\pi_{*}(1)}+\ldots+h_{n,\pi_{*}(n)}\geq\iota n\big{)}=1-o(1)\,.$
We say an algorithm $\mathcal{A}$ achieves weak partial matching, if it takes $(A,B)$ as input and outputs a family of estimators $\{h_{i,j}:1\leq i,j\leq n\}$ satisfying Items (1), (2) above and
$\mathbb{P}_{*}\big{(}h_{1,\pi_{*}(1)}+\ldots+h_{n,\pi_{*}(n)}\geq\iota n\big{)}=\Omega(1)\,.$
We point out that in the above definition $h_{i,j}$ can be thought as the estimator of $\mathbf{1}_{\{\pi_{*}(i)=j\}}$, where Item (3) implies this algorithm correctly matches a positive fraction of $\pi_{*}(i)$ with probability $1-o(1)$ (or with positive probability) and Items (1) and (2) are some regularity requirements. Our result in this section can be stated as follows:
@@
    If $(G,A,B)\sim\mathcal{G}(n,q,\rho)$ and let $\mathbb{P}_{*}$ to be the joint law of $(\pi_{*},G,A,B)$ where $\pi_{*}$ is the latent matching. Suppose that $q,\rho,D$ satisfy the assumptions in Item (1) of Theorem 3.2. Then there exists a constant $C$ such that for all $\{f_{i,j}:1\leq i,j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$, we have $\mathbb{E}_{\mathbb{P}_{*}}[f_{i,\pi_{*}(i)}]=o(1)$ for all $1\leq i\leq n$.
@@
    If $(G,A,B)\sim\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon,s)$ and let $\mathbb{P}_{*}$ to be the joint low of $(\pi_{*},G,A,B)$ where $\pi_{*}$ is the latent matching. Suppose that $\lambda,k,\epsilon,s$ and $D$ satisfy the assumptions in Item (2) of Theorem 3.2. Then there exists a constant $C$ and an event $\mathcal{E}$ such that $\mathbb{P}_{*}(\mathcal{E})=\Omega(1)$, and for all $\{f_{i,j}:1\leq i,j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ we have $\mathbb{E}_{\mathbb{P}_{*}}[f_{i,\pi_{*}(i)}\mid\mathcal{E}]=o(1)$ for all $1\leq i\leq n$.
Clearly, based on Proposition 3.3, we can deduce Theorem 3.2 via a simple Markov inequality. The rest of this section is devoted to the proof of Proposition 3.3. In the following subsections, our main focus is on proving Item (1) of Proposition 3.3. Given the similarity between the proofs of Item (1) and Item (2), for Item (2) we will provide an outline with the main differences while adapting arguments from proving Item (1) without presenting full details. [^2] After the completion of an earlier version of this draft, we realized that actually we can prove Proposition 3.3 (and thus proving Theorem 3.3) assuming only a truncated version of Conjecture 2.2 (which is weaker than Conjecture 2.5). However, this alternative proof seems to be more problem dependent so we defer it to Section C of the appendix. In the main text, we focus on proving Proposition 3.3 under the stronger assumption of Conjecture 2.5.
@@
This subsection is devoted to the proof of Item (1) in Proposition 3.3. Throughout this subsection, we will denote $\mathbb{P}_{*}$ to be the law of $(\pi_{*},G,A,B)$ where $(G,A,B)\sim\mathcal{G}(n,q;\rho)$. We will also denote $\mathbb{P}$ to be the marginal law of $(A,B)$. In addition, we assume throughout this subsection that there exists a small constant $0.010<\delta<$ 0.01 such that
@@
We first introduce some notations used in [DDL23+].
@@
We remark here that our definition of “bad” amounts to an atypically large edge density, with a carefully chosen quantitative threshold on “large”. Roughly speaking, we expect that any subgraph with size no more than $D^{2}=n^{o(1)}$ of a sparse Erdős-Rényi graph has edge-to-vertex ratio $1+o(1)$. In the definition of $\Phi$, the term $n^{1+4/D}D^{20}$ should be thought as $n^{1+o(1)}$, and $qD^{6}$ as $n^{-1+o(1)}$. The $o(1)$ terms are tuned carefully so that for a typical subgraph $H$ of a sparse Erdős-Rényi graph, $\Phi(H)$ is much bigger than $1$. The choice of $(\log n)^{-1}$ as the $\Phi$-threshold for bad graph is somewhat arbitrary, which we will only need to be vanishing as $n\to\infty$. In [DDL23+], the authors showed that one the one hand, we have $\mathbb{P}_{*}(\mathcal{E})=1-o(1)$ and thus $\operatorname{TV}(\mathbb{P}_{*},\overline{\mathbb{P}}_{*}),\operatorname{TV}(\mathbb{P},\overline{\mathbb{P}})=o(1)$; on the other hand, we have $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\mathbb{Q}}\big{)}=O_{\delta}(1)$, thus verifying the low-degree hardness for the detection problem. The first step of our proof is to generalize the result in [DDL23+], as incorporated in the following lemma.

### v4

@@ 1.4 Organization of this paper
The rest part of this paper is organized as follows: in Section 2 we state the precise framework of low-degree polynomials and how we relate it to the notion of algorithmic contiguity (see Theorem 2.4). In Section 3 we use this framework to deduce the hardness of partial matching in correlated Erdős-Rényi models and correlated SBMs, thus verifying Items (1) and (2) in Theorem 1.6 (see Theorem 3.3). In Section 4 we again use this framework to deduce the hardness of testing correlated SBMs against independent SBMs, thus verifying Item (3) in Theorem 1.6 (see Corollary 4.2). Several auxiliary proofs are moved to the appendix to ensure a smooth flow of presentation.
@@ 2 Low-degree framework and algorithmic contiguity
The low-degree polynomial framework first emerged from the works of [BHK+19, HS17, HKP+17, Hop18] and has since been refined and extended in various directions. It has found applications across a broad spectrum of problems, including detection tasks such as planted clique, planted dense subgraph, community detection, sparse PCA, and tensor PCA (see [HS17, HKP+17, Hop18, KWB22, SW22, DMW25, BKW20, DKW+22, MW25, DDL25, KMW24]), optimization problems like finding maximal independent sets in sparse random graphs [GJW20, Wein22], and constraint satisfaction problems such as random $k$-SAT [BH22]; see also the survey [KWB22]. Additionally, it is conjectured in [Hop18] that the failure of degree-$D$ polynomials implies the failure of all “robust” algorithms with running time $n^{\widetilde{O}(D)}$ (here $\widetilde{O}$ means having at most this order up to a $\operatorname{poly}\log n$ factor). However, to prevent the readers from being overly optimistic for this conjecture, we point out a recent work [BHJK25] that finds a counterexample for this conjecture. We therefore clarify that the low-degree framework is expected to be optimal for a certain, yet imprecisely defined, class of “high-dimensional” problems. Despite these important caveats, we still believe that analyzing low-degree polynomials remains highly meaningful for our setting, as it provides a benchmark for robust algorithmic performance. We refer the reader to the survey [Wein25+] for a more detailed discussion of these subtleties. In the remaining of this paper, we will focus on applying this framework in the context of high-dimensional hypothesis testing problems.
@@
We say an algorithm $\mathcal{A}$ that takes $\mathsf{Y}$ as input and outputs either 0 or $1$ achieves
  • strong detection, if the sum of type-I and type-II errors $\mathbb{Q}(\mathcal{A}(\mathsf{Y})=1)+\mathbb{P}(\mathcal{A}(\mathsf{Y})=0)$ tends to 0 as $n\to\infty$.
@@
$\mathsf{Adv}_{\leq D}(\mathbb{P},\mathbb{Q})\Big(\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\Big):=\sup_{f\in\mathcal{P}_{D}}\frac{\mathbb{E}_{\mathbb{P}}[f]}{\sqrt{\mathbb{E}_{\mathbb{Q}}[f^{2}]}}\,.$ (2.1)
@@
Note that in Conjecture 2.2 we are allowed to replace $\mathbb{P},\mathbb{Q}$ with some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ that are statistically indistinguishable with $\mathbb{P},\mathbb{Q}$. This modification enables us to avoid some situations where the low-degree advantage explodes due to some “rare events” (see [BAH+22, DDL25, DMW25] for example). Motivated by [Hop18, Hypothesis 2.1.5 and Conjecture 2.2.4] as well as the fact that low-degree polynomials capture the best known algorithms for a wide variety of statistical inference tasks, Conjecture 2.2 appears to hold for distributions of a specific form that frequently arises in high-dimensional statistics. For further discussion on what types of distributions are suitable for this framework, we refer readers to [Hop18, KWB22, Kun21, ZSWB22]. In addition, we point out that although in most applications (and in the statement of [Hop18, Hypothesis 2.1.5]) it is typically to take $\mathbb{Q}$ to be a “null” measure and $\mathbb{P}$ to be a “planted” measure (which makes (2.1) more tractable), several recent works [RSWY23, KVWX23] showed that this framework might also be applicable for many “planted-versus-planted” problems. Nevertheless, in this paper, we adopt a more conservative view and will explicitly indicate whenever $\mathbb{Q}$ is treated as a planted measure.
The framework in Conjecture 2.2 provides a useful tool for probing the computational feasibility of strong or weak detection. However, as discussed in Section 1.1, it turns out that the failure of strong detection is not enough in our cases, especially when we hope to perform some reductions between statistical models in a regime where weak (but not strong) detection is possible. Thus, in this regime, we aim to characterize a stronger framework that rules out all one-sided test. This motivates us to propose the following notion of algorithmic contiguity.
@@
For “natural” high-dimensional hypothesis testing problems between $\mathbb{P}$ and $\mathbb{Q}$, we say an algorithm $\mathcal{A}$ that takes $\mathsf{Y}$ as input and outputs either 0 or $1$ is a $\mathbb{Q}$-based one-sided test, if
@@
The goal of this section is devoted to the proof of Theorem 2.4. We first briefly explain our proof ideas. Suppose on the contrary that there is an efficient algorithm $\mathcal{A}$ that takes $\mathsf{Y}$ as input and outputs either 0 or $1$ with
$\mathbb{P}(\mathcal{A}(\mathsf{Y})=1)=\Omega(1)\mbox{ and }\mathbb{Q}(\mathcal{A}(\mathsf{Y})=0)=1-o(1)\,.$ (2.3)
$\mathbb{P}^{\prime}(\mathcal{A}(\mathsf{Y})=1)=\Omega(1)\mbox{ and }\mathbb{Q}^{\prime}(\mathcal{A}(\mathsf{Y})=0)=1-\epsilon\mbox{ for some }\epsilon=\epsilon_{n}\to 0\,.$ (2.4)
Define
$M=M_{n}=\min\big\{n,\lceil\epsilon_{n}^{-\frac{1}{2}}\rceil\big\}\to\infty\,.$ (2.5)
Definition 2.5.
  • $\overline{\mathcal{H}}_{0}$: we let $\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M}$ to be independently sampled from $\mathbb{Q}^{\prime}$.
  • $\overline{\mathcal{H}}_{1}$: we first sample $\kappa\in\{1,\ldots,M\}$ uniformly at random, and (conditioned on the value of $\kappa$) we let $\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M}$ are independent samples with $\mathsf{Y}_{\kappa}$ generated from $\mathbb{P}^{\prime}$ and $\{\mathsf{Y}_{j}:j\neq\kappa\}$ generated from $\mathbb{Q}^{\prime}$.
In addition, denote $\overline{\mathbb{P}}$ and $\overline{\mathbb{Q}}$ to be the law of $(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M})$ under $\overline{\mathcal{H}}_{1}$ and $\overline{\mathcal{H}}_{0}$, respectively.
Now assuming that (2.4) holds, we see that
[…]
$\overline{\mathbb{P}}\Big(\big(\mathcal{A}(\mathsf{Y}_{1}),\ldots,\mathcal{A}(\mathsf{Y}_{M})\big)\neq(0,\ldots,0)\Big)\geq\Omega(1)\,.$ (2.7)
Lemma 2.6.
Proof.
Note that
@@
$\frac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\overline{\mathbb{Q}}}(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M})=\frac{1}{M}\sum_{i=1}^{M}\frac{\mathrm{d}\overline{\mathbb{P}}(\cdot\mid\kappa=i)}{\mathrm{d}\overline{\mathbb{Q}}}(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M})=\frac{1}{M}\sum_{i=1}^{M}\frac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}(\mathsf{Y}_{i})\,.$ (2.8)
In addition, recall (2.1). Define
[…]
$\mathsf{Adv}_{\leq D}\big(\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big)=\Bigg(\sum_{\alpha\in\Lambda}\mathbb{E}_{\mathbb{P}^{\prime}}[f_{\alpha}(\mathsf{Y})]^{2}\Bigg)^{\frac{1}{2}}\,.$ (2.9)
For any $f\in\mathcal{N}_{D}^{\perp}$, it can be uniquely expressed as
$f=\sum_{\alpha\in\Lambda}C_{\alpha}f_{\alpha}\,$,
with equality holds if and only if $C_{\alpha}\propto\mathbb{E}_{\mathbb{P}^{\prime}}[f_{\alpha}]$. This yields (2.9). Now, note that $\overline{\mathbb{Q}}=(\mathbb{Q}^{\prime})^{\otimes M}$ is a product measure of $\mathbb{Q}^{\prime}$, there is a natural standard orthogonal basis under $\overline{\mathbb{Q}}$, given by
Thus, similarly as in (2.9), we see that
$\Big(\mathsf{Adv}_{\leq D}\big(\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\overline{\mathbb{Q}}}\big)\Big)^{2}=\sum_{\begin{subarray}{c}(\alpha_{1},\ldots,\alpha_{M}):\alpha_{i}\in\Lambda\\ \sum_{i=1}^{M}\operatorname{deg}(f_{\alpha_{i}})\leq D\end{subarray}}\mathbb{E}_{\overline{\mathbb{P}}}\Big[\prod_{i=1}^{M}f_{\alpha_{i}}(\mathsf{Y}_{i})\Big]^{2}\,.$ (2.10)
In addition, using (2.8), we see from direct calculation that
[…]
$=\begin{cases}1\,,&(\alpha_{1},\ldots,\alpha_{M})=(0,\ldots,0)\,;\\ \frac{1}{M}\mathbb{E}_{\mathbb{P}^{\prime}}\big[f_{\alpha_{j}}(\mathsf{Y}_{j})\big]\,,&(\alpha_{1},\ldots,\alpha_{M})=(0,\ldots,0,\alpha_{j},0,\ldots,0)\,;\\ 0\,,&\mbox{otherwise}\,.\end{cases}$ (2.11)
Plugging (2.11) into (2.10), we get that
(2.10) $=1+\sum_{i=1}^{M}\sum_{\alpha_{i}\in\Lambda\setminus\{0\}}\Big(\frac{1}{M}\mathbb{E}_{\mathbb{P}^{\prime}}\big[f_{\alpha_{i}}(\mathsf{Y}_{i})\big]\Big)^{2}$
[…]
where in the second equality we use (2.9) and the assumption that $\mathsf{Adv}_{\leq D}\big(\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big)=O(1)$. This completes our proof. ∎
We can now finish the proof of Theorem 2.4.
Proof of Theorem 2.4.
Suppose on the contrary that there is an efficient algorithm $\mathcal{A}$ satisfying (2.3). Consider the hypothesis testing problem stated in Definition 2.5. Using (2.6) and (2.7), we see that there is an efficient algorithm that achieves weak detection between $\overline{\mathbb{P}}$ and $\overline{\mathbb{Q}}$, which contradicts with Lemma 2.6 and Item (2) in Conjecture 2.2. ∎
@@ 3 Partial recovery in correlated random graphs
For two elements $\pi,\pi^{\prime}\in\mathfrak{S}_{n}$, define
$\mathsf{OV}(\pi,\pi^{\prime})=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{\pi(i)=\pi^{\prime}(i)\}}\,.$ (3.1)
Given a sample $(A,B)$ from the law a pair of correlated random graphs in Definition 1.1 (we denote this law of $\mathbb{P}_{*}$). We say an algorithm $\mathcal{A}$ achieves strong partial matching, if it takes $(A,B)$ as input and outputs an estimator $\widehat{\pi}=\widehat{\pi}(A,B)\in\mathfrak{S}_{n}$ such that there exists a fixed constant $\iota>0$ with
$\mathbb{P}_{*}\big(\mathsf{OV}(\widehat{\pi},\pi_{*})\geq\iota\big)=1-o(1)\,.$ (3.2)
We say an algorithm $\mathcal{A}$ achieves weak partial matching, if it takes $(A,B)$ as input and outputs an estimator $\widehat{\pi}=\widehat{\pi}(A,B)\in\mathfrak{S}_{n}$ such that there exists a fixed constant $\iota>0$ with
$\mathbb{P}_{*}\big(\mathsf{OV}(\widehat{\pi},\pi_{*})\geq\iota\big)=\Omega(1)\,.$ (3.3)
Remark 3.2.
Given any estimator $\widehat{\pi}(A,B)\in\mathfrak{S}_{n}$, by defining $h_{i,j}=\mathbf{1}_{\widehat{\pi}(i)=j}$ we get a family of estimators $\{h_{i,j}:1\leq i,j\leq n\}$ such that
@@
    $h_{i,1}+\ldots+h_{i,n}=1$ for all $1\leq i\leq n$ a.s. under $\mathbb{P}_{*}$.
In addition, suppose $\widehat{\pi}(A,B)$ achieves strong partial recovery, it is easy to check that
@@
    $\mathbb{P}_{*}(h_{1,\pi_{*}(1)}+\ldots+h_{n,\pi_{*}(n)}\geq\iota n)=1-o(1)$.
And suppose $\widehat{\pi}(A,B)$ achieves weak partial recovery, it is easy to check that
  1. (3’)
    $\mathbb{P}_{*}(h_{1,\pi_{*}(1)}+\ldots+h_{n,\pi_{*}(n)}\geq\iota n)=\Omega(1)$.
In the rest part of this section, we will use the estimator $\widehat{\pi}$ and the family of estimators $\{h_{i,j}:1\leq i,j\leq n\}$ interchangeably.
Our result in this section can be stated as follows:
@@
    If $(G,A,B)\sim\mathcal{G}(n,q,\rho)$ and let $\mathbb{P}_{*}$ to be the joint law of $(\pi_{*},G,A,B)$ where $\pi_{*}$ is the latent matching. Suppose that $q,\rho,D$ satisfy the assumptions in Item (1) of Theorem 3.3. Then there exists a constant $C$ such that for all $\{h_{i,j}(A,B):1\leq i,j\leq n\}$ that satisfies Items (1),(2) in Remark 3.2 and can be computed in time $n^{D/(\log n)^{C}}$, we have $\mathbb{E}_{\mathbb{P}_{*}}[h_{i,\pi_{*}(i)}]=o(1)$ for all $1\leq i\leq n$.
@@
    If $(G,A,B)\sim\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon,s)$ and let $\mathbb{P}_{*}$ to be the joint law of $(\pi_{*},G,A,B)$ where $\pi_{*}$ is the latent matching. Suppose that $\lambda,k,\epsilon,s$ and $D$ satisfy the assumptions in Item (2) of Theorem 3.3. Then there exists a constant $C$ and an event $\mathcal{E}$ such that $\mathbb{P}_{*}(\mathcal{E})=\Omega(1)$, and for all for all $\{h_{i,j}(A,B):1\leq i,j\leq n\}$ that satisfies Items (1),(2) in Remark 3.2 and can be computed in time $n^{D/(\log n)^{C}}$, we have $\mathbb{E}_{\mathbb{P}_{*}}[h_{i,\pi_{*}(i)}\mid\mathcal{E}]=o(1)$ for all $1\leq i\leq n$.
Clearly, based on Proposition 3.4, we can deduce Theorem 3.3 via a simple Markov inequality. The rest of this section is devoted to the proof of Proposition 3.4. In the following subsections, our main focus is on proving Item (1) of Proposition 3.4. Given the similarity between the proofs of Item (1) and Item (2), for Item (2) we will provide an outline with the main differences while adapting arguments from proving Item (1) without presenting full details.
@@
This subsection is devoted to the proof of Item (1) in Proposition 3.4. Throughout this subsection, we will denote $\mathbb{P}_{*}$ to be the law of $(\pi_{*},G,A,B)$ where $(G,A,B)\sim\mathcal{G}(n,q;\rho)$. We will also denote $\mathbb{P}$ to be the marginal law of $(A,B)$. In addition, we assume throughout this subsection that there exists a small constant $0<\delta<0.01$ such that
@@
We first introduce some notations used in [DDL25].
@@
We remark here that our definition of “bad” amounts to an atypically large edge density, with a carefully chosen quantitative threshold on “large”. Roughly speaking, we expect that any subgraph with size no more than $D^{2}=n^{o(1)}$ of a sparse Erdős-Rényi graph has edge-to-vertex ratio $1+o(1)$. In the definition of $\Phi$, the term $n^{1+4/D}D^{20}$ should be thought as $n^{1+o(1)}$, and $qD^{6}$ as $n^{-1+o(1)}$. The $o(1)$ terms are tuned carefully so that for a typical subgraph $H$ of a sparse Erdős-Rényi graph, $\Phi(H)$ is much bigger than $1$. The choice of $(\log n)^{-1}$ as the $\Phi$-threshold for bad graph is somewhat arbitrary, which we will only need to be vanishing as $n\to\infty$. In [DDL25], the authors showed that one the one hand, we have $\mathbb{P}_{*}(\mathcal{E})=1-o(1)$ and thus $\operatorname{TV}(\mathbb{P}_{*},\overline{\mathbb{P}}_{*}),\operatorname{TV}(\mathbb{P},\overline{\mathbb{P}})=o(1)$; on the other hand, we have $\mathsf{Adv}_{\leq D}\big(\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\mathbb{Q}}\big)=O_{\delta}(1)$, thus verifying the low-degree hardness for the detection problem. The first step of our proof is to generalize the result in [DDL25], as incorporated in the following lemma.